How fast does an 80 kg person accelerate when pulled with the force given below in Newtons? a. 1 b. 2 c. 5 d. 10 e. 20 f. 50 g. 100 h. 200 i. 500 j. 1000 k. 2000
Answer #1

m= 80kg. a> F=1N, as pro = Ta=0.0125 m/s2 l a 0.01 by F=an , a = 7 th = 2/50 19-01025 m[8] c) &F=5N, a= 2 10 Ta=0:06a5 in 182g) F=100N , as fo = 100 I a= 1.25m/s² S n> F= 2001, as to the a=2.5 m/s2 i> F=1000N , a = Bat 1000 80 a=12-5 m/s² j> F= 2000N